A series R-L circuit is connected to 110V ac source .If voltage across resistor is 85V find power dissipated in circuit


1
Expert's answer
2021-05-31T17:09:32-0400

Since the voltages across the inductor and the resistor are "90^0"out of phase, we obtain the voltage across the inductor in the phase domain as follows: If the voltage across the resistor is"V_1", the voltage across the inductor will be"jV_2".

Assume the voltage of the

source is

"V_3"

Applying KVL law


"V_1+jV_2=V_1cos(wt)+V_2sin(wt)=V_3cos(wt-\theta)=85+V_2sin(wt)=110cos(wt-\theta)"

"110=\sqrt{85^2+V_2^2}"

"V_2=69.82V"

For example

"P_{ave}=\frac{1}{2}\frac{V^2}{R}"

Where V=85V

Assume "R=20 \Omega"

Put value

"P_{av}=\frac{1}{2}\frac{85^2}{20}=180.6 W"